A toll bridge charges $4 for a car and $ 6 for a truck.
One day 200 of these vehicles cross the bridge and paid a total of $ 860.
How many we're trucks?

Answer by checkley79(3341)

4C+6T=860
C+T=200
C=200-T
4(200-T)+6T=860
800-4T+6T=860
2T=860-800
2T=60
T=60/2
T=30 TRUCKS.
C+30=200
C=200-30
C=170 CARS.
PROOF:
4*170+6*30=860
680+180=860
860=860